If I invest 3 million dollars, and I believe I will have sales of 100k in the first year and after that sales will increase by 25% per year for the next 3 years. After that sales will grow at a constant 3% rate forever. Market risk premium of 4.4%. 1 year tbill yield curve of 2.6%. Beta is .52. Should I make this investment? Why or why not?
Answer #1

Cost of Equity =Risk free Rate+Beta*Market Risk Premium =2.6%+0.52*4.4% =4.888%
Initial Investment =3,000,000
Sales in one year =100000
Sales in 2 years =100000*1.25 =125000
Sales in 3 years =100000*1.25^2 =156250
Sales in 4 years =100000*1.25^3= 195312.50
Terminal Value =195312.50/(4.888%-3%) =10,344,941.74
Value of this investment =100000/(1+4.888%)+125000/(1+4.888%)^2+156250/(1+4.888%)^3+10344941.74/(1+4.888%)^3-3,000,000
=9,27309,375-3,000,000 =4731534.02
Yes this investment should be undertaken. because value is positive.